OpenStudy (anonymous):

I give medalsssss Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by five times cosine of pi divided by seven plus sine of pi divided by five times sine of pi divided by seven.

OpenStudy (anonymous):

This is how far ive gotten cos(pi/5-pi/7)+sin(pi/5-pi/7) cos(7pi/35-5pi/35)+sin(7pi/35-5pi/35) cos(2pi/35)+sin(2pi/35)

OpenStudy (anonymous):

@thomaster

OpenStudy (mathmate):

Is it: \(cos\frac{\pi}{5}*cos\frac{\pi}{7}+sin\frac{\pi}{7}sin\frac{\pi}{5}\) If the above expression is correct, then you can use the difference formula: \(cos(x-y)=cos(x)cos(y)+sin(x)sin(y)\)

OpenStudy (anonymous):

isnt that what i did?

OpenStudy (mathmate):

I do not know what you solved because the problem you solved is different from what i posted. So please confirm what the problem is before we can figure out solutions.

OpenStudy (anonymous):

cos(pi/5)cos(pi/7)+sin(pi/5)sin(pi/7)

OpenStudy (mathmate):

ok, then the identity I suggested would help you. You were close, but for a different reason.

OpenStudy (anonymous):

i dont really understand what the question is asking...

OpenStudy (mathmate):

The question wants you to recognize that the expression is either the sin, cos or tan function of a single angle, like cos(2pi/35), or sin(pi/24), etc. If you examine the right-hand side of the identity I proposed, you will be able to convert the complicated expression into a single function of a single angle.

OpenStudy (anonymous):

ok thanks!

OpenStudy (mathmate):

You're welcome! :)